OpenStudy (anonymous):

Find the limit of the function algebraically. limit as x approaches negative six of quantity x squared minus thirty six divided by quantity x plus six.

OpenStudy (anonymous):

I just need to know if I can use direct substitution right away or if there are any step I need to do before that

OpenStudy (anonymous):

can you try writing it with the equation?

OpenStudy (anonymous):

sorry its hard for me to visualize the problem purely with words

OpenStudy (anonymous):

http://prntscr.com/80cyqz

OpenStudy (anonymous):

\[\lim_{x \rightarrow -6}\frac{ x^2-36 }{ x+6 } = \frac{ (x-6)(x+6) }{ x+6 }\]now try to continue from here

OpenStudy (anonymous):

x-6

OpenStudy (anonymous):

yea

OpenStudy (loser66):

just small mistake \(\lim_{x \rightarrow -6}\frac{ x^2-36 }{ x+6 } =lim_{x\rightarrow -6}\frac{ (x-6)(x+6) }{ x+6 }=\color{red}{lim_{x\rightarrow -6}}x-6\) Then replace x =-6 in It is small but very important because lim (something) = value, not = x -6

OpenStudy (anonymous):

thanks